# Rethinking Weight-Averaged Model-merging

### v2

@@ 6.3 Model Predictions and Weight Magnitudes
Specifically, we visualize the merging of 10 models (merging of 2 to 7 models can be found in the supplementary materials) for model-merging and ensemble towards different magnitude factors, as shown in Fig. 8(b). In the figure, “Model Perf” denotes the average of individual model performance without magnifying the weights (that is why it is a constant value), and “Model ×$\times M$ Perf” represents the model with weight magnification. From the trends in the figures, we can observe the model-merging models are more robust than model ens models, including ensemble on logits and features. But surprisingly, feature ensembles are more robust than logits ensembles across different configurations. Feature ensemble models can even outperform model merging when weight magnification is not large (e.g., ×$\times 50$). This shows the feature ensemble can have some resistance towards the change of weight magnitudes and variances. The “Ens Features” model uses the merge of the last fully connected (FC) layers from all merged models for classification. In order to mitigate the impact of the merged FC layer, we introduce a variant of the model, “Ens Features*”, by directly adopting the FC layer from the 1st model. We can observe the “Ens Features” and “Ens Features*” models perform similarly.
We further visualize the gaps between weight-averaged model-merging and ensemble in Fig. 8(c). When more models are taken into account, the gaps become larger, but there is a peak point for the gap (×$\times 100$ for Cifar100). The same experiments to visualize the merging of 2 to 10 models for Cifar10 with different magnitude factors, as well as the gaps between weight-averaged model-merging and ensemble, are shown in the supplementary materials. Also, a similar phenomenon is shown on Cifar10, but with different peaks of gaps w.r.t. the magnitudes. Moreover, we conduct the weight magnitude experiment on a relatively large dataset, Tiny ImageNet, in Tab. 3. From the table, we can observe that model merging can still mitigate the weight magnifying problem to a certain extent. It can also outperform logit ensembles, but can not improve over ensemble on features and is not as good as on Cifar100 (e.g. 4.81% to 6.86% accuracy for 10 models merging). This may be caused by the complexity of the Tiny ImageNet task and as we discussed in the Sec. 5.3, the model-merging can work to some extent as it performs similar operations as regularization (e.g., L1 or L2 regularization). It will mitigate the weight magnitude changing issues, but with the cost of losing model expressiveness.
Table 3: Model-merging to address the weight magnitude issue for VGG19 with different model numbers on Tiny ImageNet dataset with the magnitude factor ×$\times 100$. This approach works for smaller data (Cifar10/Cifar100) as it plays a regularization role. However, it does not work for larger datasets such as Tiny ImageNet, because it sacrifices the expressiveness of the models.
@@ 7 Conclusion
In this paper, we thoroughly examined the weight-averaged model-merging approach through the lens of template matching by visualizing weight patterns across various datasets. Our comparison of weight and feature averaging within ensemble and merging methods revealed notable differences in their effects on different model architectures and datasets, offering insights into when and why each approach is beneficial. Furthermore, we demonstrate that model merging techniques tend to yield more stable predictions than model ensembles, especially when subjected to variations in parameter magnitudes, thus enhancing robustness and generalization. These findings clarify the interpretability and key aspects of model-merging, providing a deeper understanding of its application across diverse contexts. We anticipate that these insights will stimulate further research into model merging strategies and support the development of more interpretable and robust model-averaging methodologies.
@@ 8 Experimental Settings
For the experimental settings of model weight patterns. We train one single linear layer on both Cifar100 [12] and Tiny ImageNet [13] datasets with momentum SGD as the optimizer. The initial learning rater is set to 0.01 and decreases every 20 epochs by 0.1. While visualization, the weights are wrapped back to the image size.
@@
In terms of averaging on weights versus averaging on features, we consider Uniform Soups and Greedy Soups [33] for weight-averaged model-merging. We design the experiments of this section from two different angles: model-wise examination and dataset-wise examination of the model-merging and model ensemble performances. For the models considered, we have ResNet [8] (ResNet18 to ResNet152), DenseNet [9] (DenseNet121 to DenseNet201), VGG [22] (VGG11 to VGG19, including with and without batch normalization), ViT [24, 6] and DeiT [27]. For the dataset-wise examination, we conduct experiments on Cifar10/Cifar100, PathMNIST/DermaMNIST [37, 38], Celeba [15], Tiny ImageNet and ChestXRay [31, 16]. The information of these datasets can be find in Tab. 4. Specially, ChestXRay dataset has 14 labels and each of them is a binary classification task. We demonstrate the average Area Under the Receiver Operating Characteristic Curve (AUROC) across 14 labels as the model performance for ChestXRay. For model ensemble, we consider ensemble on model logits (before entering softmax or sigmoid) and ensemble on model features (before the final classification layer). Besides these two model ensemble techniques, whenever a model is considered to be merged into a greedy soup model, we ensemble it using either the Greedy Ensemble Logits model (Grd Ens Lgt) or Greedy Ensemble Features model (Grd Ens Fts). In order to exclude the impact of adopting the last layer from any single model, the “Ens Features” model uses the merging of the last fully connected (FC) layers from all merged models for classification.
@@
For the experiments of model predictions and weight magnitudes, we conduct experiments on Cifar10, Cifar100 and Tiny ImageNet datasets. To increase the magnitude and the variance of model weights, we magnify the model weights by a constant factor. Formally, when each element in a matrix $A$ is scaled by a constant factor $c$, the variance of the resulting matrix elements increases proportionally to the square of this factor. For example, in the case where $c=100$, the variance of the elements of matrix $A$ becomes $100^{2}\times\text{Var}(A)=10,000\times\text{Var}(A)$.
@@
For all experiments in the paper, to keep fair comparisons, we adopt the official splits for each dataset. Also, we uniformly train 10 models and perform the model-merging and ensemble under exactly same conditions (learning rates, step strategies and optimizers, etc).
@@ 9 More Visualization of Model Weight Patterns
Similar to the visualization in the main paper, we visualize the average of all images Fig. 9 and linear classifiers for all classes Fig. 10 within each class (all 100 classes) of Cifar100 dataset. We can easily and clearly observe the strong correlation between the average of images in a class and the learned weights.
@@
More visualization of the convolutional kernels on Cifar100 within deep learning models, e.g. Fig. 12 for ResNet50 and Fig. 13 for VGG19. For ResNet50 kernel visualization, we can observe differences or similarities between each learned patterns. For instance, Kernel 5 and kernel 9 are similar, which are responsible for detecting similar features and thus yield a higher measurement (e.g. inner product) between the two kernels; while Kernel 9 is very different with Kernel 45.The patterns inside ResNet50 is easier to be observed than VGG19, as ResNet50 has larger kernel size for the 1st layer.
@@
For Tiny ImageNet, we have Fig. 14 and Fig. 16 for the average of all images within 200 classes in the dataset; also, Fig. 15 and Fig. 17 for the linear classifier visualization of all 200 classes in the dataset. We can observe the correlations between averaged images and the learned patterns.
@@
Fig. 22 shows the performance gaps between ensembling on logits with greedy soups. From it, we can observe similar phenomenon with the gaps between ensembling on logits with uniform soups, with more models merged, the larger gaps between ensemble on logits with greedy soups will be. But differently, the greedy soups break the going down trend with uniform soups as described in the main paper.
@@
We illustrate the ViT model (DeiT Tiny) on another dataset — Tiny ImageNet in Tab. 10. We can also see the more models to be merged, the poorer results from uniform soups can get. Also, we can observe the greedy soups, greedy ensemble on logits and greedy ensemble on features are not changed, it further demonstrates the model-merging between 2 or more models will not contribute to the performance.
@@ 11 More Experimental Results of Model Magnitude Changes
The 2 models, 3 models, 5 models and 7 models accuracy vs. magnification factor on Cifar100 are shown via Fig. 23, Fig. 24, Fig. 25 and Fig. 26, respectively.
@@
Similar to Cifar100, the 2 models, 3 models, 5 models, 7 models and 10 models accuracy vs. magnification factor on Cifar10 are shown via Fig. 27, Fig. 28, Fig. 29, Fig. 30 and Fig. 31, respectively. Moreover, we also visualize the uniform soup and logits ensemble accuracy vs. magnification factor across different model numbers on the Cifar10 dataset in Fig. 32. We can observe similar phenomenons. However, on the Cifar10, the gaps have different a peak point, which is ×$\times 110$ (for Cifar100, it is ×$\times 100$).

### v3

@@ 6.3 Model Predictions and Weight Magnitudes
Moreover, we visualize 10 models (merging of 2 to 7 models can be found in the supplementary materials) merging and ensembling towards different magnitude factors, as shown in Fig. 8(b). In the figure, “Model Perf” denotes the average of individual model performance without magnifying the weights (that is why it is a constant value), and “Model ×$\times M$ Perf” represents the model with different magnification. From the trends in the figures, we can observe the merged models are more robust than ensembled ones (with the factor grows, the line denoting ensembling goes down quickly), including ensemble on logits and features. But interestingly, feature ensembling is more robust than logits ensembling across different configurations. Feature ensembling can even outperform model-merging when weight magnification is not large (e.g., ×$\times 50$). This shows the feature ensembling can have some resistance towards the change of weight magnitudes. The “Ens Features” performs ensemble on multiple model features, while merging all last fully connected (FC) layers to get rid of the influence of FC layer. In order to mitigate the impact of the merged FC layer, we introduce a variant of the model, “Ens Features*”, by directly adopting the FC layer from the 1st model. We can observe the “Ens Features” and “Ens Features*” models perform similarly. It shows the low sensitivity of the merging of last layers.
We further visualize the gaps between model-merging and ensembling in Fig. 8(c). When more models are taken into account, the gaps become larger, but there is a peak point for the gap (×$\times 100$ is the peak for Cifar100). The experiments of merging/ensembling 2 to 10 models with different magnitude factors, as well as on different datasets and the gaps between merging and ensembling, are shown in the supplementary materials. A similar phenomenon is shown on different datasets, but with different peaks. Moreover, we conduct the weight magnitude experiment on a relatively large dataset, Tiny ImageNet, in Tab. 3. From the table, we can observe that model-merging can still mitigate the weight magnifying problem to a certain extent. It can also outperform logit ensembling, and cannot outperform feature ensembling. But generally it cannot work very well as on Cifar100. It may be caused by the complexity of the Tiny ImageNet and as we discussed in the Sec. 5.3, the model-merging can work to some extent as it can act similarly as regularization. It can mitigate the weight magnitude increment issues, but with the cost of losing model expressiveness.
Table 3: Model-merging to address the weight magnitude issue for VGG19 with different model numbers on Tiny ImageNet dataset with the magnitude factor ×$\times 100$. This approach works for smaller data (Cifar10/Cifar100) as it has a regularization effect. However, it does not work for larger datasets such as Tiny ImageNet, because it sacrifices the expressiveness of the models.
@@ 7 Conclusion
In this paper, we thoroughly examined the weight-averaged model-merging approach through the lens of template matching by visualizing weight patterns across various datasets. Our comparison of weight and feature averaging within ensembling and merging methods revealed notable differences in their effects on different architectures and datasets, offering insights into when and why each approach is beneficial. Furthermore, we demonstrate that model-merging techniques tend to yield more stable predictions than model ensembles, especially when subjected to variations in parameter magnitudes, thus enhancing robustness and generalization. These findings clarify the interpretability and key aspects of model-merging, providing a deeper understanding of its application across diverse contexts. We anticipate that these insights will stimulate further research into model-merging strategies and support the development of more interpretable and robust model-averaging methodologies.
@@ 8 Experimental Settings
For the experimental settings of model weight patterns. We train one single linear layer on both Cifar100 [14] and Tiny ImageNet [15] datasets with momentum SGD as the optimizer. The initial learning rater is set to 0.01 and decreases every 20 epochs by 0.1. While visualization, the weights are wrapped back to the image size as described in the main paper (i.e. 100 $\times$ 3 $\times$ 32 $\times$ 32 for Cifar100).
@@
In terms of averaging on weights versus averaging on features, we consider Uniform Soups and Greedy Soups [35] for weight-averaged model-merging, which is the focus of our paper. We design the experiments of this section from two different angles: model-wise examination and dataset-wise examination of the model-merging and model ensemble performances. For the models considered, we have ResNet [9] (ResNet18 to ResNet152), DenseNet [10] (DenseNet121 to DenseNet201), VGG [24] (VGG11 to VGG19, including with and without batch normalization), ViT [26, 5] and DeiT [29]. For the dataset-wise examination, we conduct experiments on Cifar10/Cifar100, PathMNIST/DermaMNIST [39, 40], Celeba [17], Tiny ImageNet and ChestXRay [33, 18]. The information of these datasets can be find in Tab. 4. Specially, ChestXRay-14 is a chest X-ray imaging dataset designed for multi-label classification. It consists of 112,120 frontal chest X-ray images from 30,805 unique patients. The dataset includes annotations for 14 different lung disease categories. We use the official dataset split for training and evaluation. We demonstrate the average Area Under the Receiver Operating Characteristic Curve (AUROC) across 14 labels as the model performance for ChestXRay.
For model ensembling, we consider to ensemble on model logits (before entering softmax or sigmoid) and ensemble on model features (before the final classification layer). Besides these two model ensembling techniques, whenever a model is considered to be merged into a greedy soup model, we also ensemble it with either we called the Greedy Ensemble Logits model (Grd Ens Lgt) or Greedy Ensemble Features model (Grd Ens Fts). In order to exclude the impact of adopting the last layer from any single model, the “Ens Features” model uses the merging of the last fully connected (FC) layers from all merged models for classification.
@@
For the experiments of model predictions and weight magnitudes, we conduct experiments on Cifar10, Cifar100 and Tiny ImageNet datasets. To increase the magnitude and the variance of model weights, we magnify the model weights by a magnitude factor. Formally, when each element in a matrix $A$ is scaled by a factor $c$, the variance of the resulting matrix elements increases proportionally to the square of this factor. For example, in the case where $c=100$, the variance of the elements of matrix $A$ becomes $100^{2}\times\text{Var}(A)=10,000\times\text{Var}(A)$.
@@
For all experiments in the paper, to keep fair comparisons, we adopt the official splits for each dataset. If they provide publicly available dataset with Python interface, we directly call the interface to import the data. Also, we uniformly train 10 models and perform the model-merging and ensembling under exactly same conditions (learning rates, step strategies and optimizers, model numbers, etc).
9 Theoretical Proofs
For Eq. 14 in the paper, we aim to prove that:
$\frac{1}{4}(\sigma_{1}^{2}+\sigma_{2}^{2})\leq\max(\sigma_{1}^{2},\sigma_{2}^{2})$
Assume without loss of generality that $\sigma_{1}^{2}\leq\sigma_{2}^{2}$ (the argument is symmetric if $\sigma_{2}^{2}\leq\sigma_{1}^{2}$).
Given that $\max(\sigma_{1}^{2},\sigma_{2}^{2})=\sigma_{2}^{2}$, the inequality becomes:
$\frac{1}{4}(\sigma_{1}^{2}+\sigma_{2}^{2})\leq\sigma_{2}^{2}$
Now, observe the following:
$\frac{1}{4}(\sigma_{1}^{2}+\sigma_{2}^{2})\leq\frac{1}{4}(2\sigma_{2}^{2})=\frac{\sigma_{2}^{2}}{2}$
Since $\frac{\sigma_{2}^{2}}{2}\leq\sigma_{2}^{2}$, we conclude that:
$\frac{1}{4}(\sigma_{1}^{2}+\sigma_{2}^{2})\leq\sigma_{2}^{2}=\max(\sigma_{1}^{2},\sigma_{2}^{2})$
Thus, the inequality holds true.
For the Property 1 and Theorem 1, according to [34], we have:
Lemma 1: Given a fully-connected network, its weight matrix $\mathbf{W}^{(m)}$ satisfies, for any $\tau>0$,
$\small\norm{\mathbf{W}^{(m)}}_{2}\leq\sqrt{N}+C_{s}K_{s}^{2}(\sqrt{N}+\tau),$
with a probability of at least $1-2e^{-\tau^{2}}$, where $C_{s}$ is a universal constant and $K_{s}=\max_{i}\|\mathbf{W}_{i}^{(m)}\|_{2}$.
For Property 1:
Proof.
We denote $s_{1}(A)$ as the maximum singular value of $A$ and set $\lambda=\sqrt{N}+C_{s}K_{s}^{2}(\sqrt{N}+\tau)$ as the upper bound of the maximum singular value of all weight matrices by Lemma 1. By Lipschiz property of the activation function $\phi$ and $\norm{A\mathbf{x}}=s_{1}(A)\norm{\mathbf{x}}$,
$\norm{\mathbf{y}^{(m)}}_{2}=$ $\norm{\phi(\mathbf{W}^{(m)}\mathbf{y}_{k}^{(m-1)})}_{2}$
$=$ $\norm{\phi(\mathbf{W}^{(m)}\mathbf{y}_{k}^{(m-1)})-\phi(0)}_{2}$
$\leq$ $L\norm{\mathbf{W}^{(m)}\mathbf{y}^{(m-1)}-0}_{2}$
$=$ $L\norm{\mathbf{W}^{(m)}\mathbf{y}^{(m-1)}}_{2}$
$=$ $Ls_{1}(\mathbf{W}^{(m)})\norm{\mathbf{y}^{(m-1)}}_{2}$
$\leq$ $L\lambda\norm{\mathbf{y}^{(m-1)}}_{2}$
$\leq$ $(L\lambda)^{2}\norm{\mathbf{y}^{(m-2)}}_{2}$
$\cdots$
$\leq$ $(L\lambda)^{m}\norm{\mathbf{x}}_{2}.$
Then we calculate the probability that this inequality holds by Lemma 1,
$\mathbb{P}\left(\norm{\mathbf{y}^{(m)}}_{2}\leq(L\lambda)^{m}\norm{\mathbf{x}}_{2}\right)$
$=$ $\prod_{k=1}^{m}\mathbb{P}\left(\norm{\mathbf{W}^{(k)}\mathbf{y}^{(k-1)}}_{2}\leq s_{1}(\mathbf{W}^{(k)})\norm{\mathbf{y}^{(k-1)}}_{2}\right)$
$\geq$ $(1-2e^{-\tau^{2}})^{m}.$
For Theorem 1:
Proof.
For the first hidden layer, we have a linear transformation for one neuron:
$\mathbf{W}^{(1)}_{i}\mathbf{x}+b^{(1)}_{i}.$
Given that all weights and biases are i.i.d, we get:
$\mathbb{V}(\mathbf{W}^{(1)}_{i}\mathbf{x}+b^{(1)}_{i})$ $=\mathbb{V}\left(\sum^{N^{(1)}}_{j=1}\mathbf{W}^{(1)}_{i,j}x_{j}+b^{(1)}_{i}\right)$
$=\mathbb{V}\left(\sum^{N^{(1)}}_{j=1}\mathbf{W}^{(1)}_{i,j}x_{j}\right)+\mathbb{V}(b^{(1)}_{i})$
$=\sum^{N^{(1)}}_{j=1}x_{j}^{2}\mathbb{V}(\mathbf{W}^{(1)}_{i,j})+\mathbb{V}(b^{(1)}_{i})$
$=(\sigma^{(1)}_{w})^{2}\sum^{N^{(1)}}_{j=1}x_{j}^{2}+(\sigma^{(1)}_{b})^{2}$
$=(\sigma^{(1)}_{w})^{2}\|\mathbf{x}\|_{2}^{2}+(\sigma^{(1)}_{b})^{2}.$
By the Lipschitz property of the activation function, we have:
$f^{(1)}_{i}=\phi(\mathbf{W}^{(1)}_{i}\mathbf{x}+b^{(1)}_{i})\leq L(\mathbf{W}^{(1)}_{i}\mathbf{x}+b^{(1)}_{i}).$
Then,
$\mathbb{V}(f^{(1)}_{i})$ $\leq\mathbb{V}[L(\mathbf{W}^{(1)}_{i}\mathbf{x}+b^{(1)}_{i})]$
$\leq L^{2}\mathbb{V}(\mathbf{W}^{(1)}_{i}\mathbf{x}+b^{(1)}_{i})$
$\leq L^{2}\left[(\sigma^{(1)}_{w})^{2}\|\mathbf{x}\|_{2}^{2}+(\sigma^{(1)}_{b})^{2}\right].$
Hence,
$\mathbb{V}(f^{(1)})$ $=\sum^{N}_{j=1}\mathbb{V}(f^{(1)}_{i})$
$\leq L^{2}N\left[(\sigma^{(1)}_{w})^{2}\|\mathbf{x}\|_{2}^{2}+(\sigma^{(1)}_{b})^{2}\right].$
Similarly, for other layers:
$\mathbb{V}(f^{(m)})\leq L^{2}N^{(m-1)}\left[(\sigma^{(m)}_{w})^{2}\|f^{\prime(m-1)}\|_{2}^{2}+(\sigma^{(m)}_{b})^{2}\right].$
This result follows by iteratively applying the inequalities, concluding the proof. ∎
@@ 10 More Visualization of Model Weight Patterns
Similar to the visualization in the main paper, we visualize the average of all images Fig. 9 and linear classifiers for all classes Fig. 10 within each class (all 100 classes) of Cifar100 dataset. We can clearly observe the strong correlation between the average of images per class and its learned weights.
@@
More visualization of the convolutional kernels on Cifar100 within deep learning models, e.g. Fig. 12 for ResNet50 and Fig. 13 for VGG19. For ResNet50 kernel visualization, we can observe differences or similarities between each learned patterns. For instance, Kernel 5 and kernel 9 are similar in terms of textures and grayscale colors, which are responsible for detecting similar features and thus yield a similar activations (e.g. inner product) when interact with inputs; while Kernel 9 is very different with Kernel 45. The patterns inside ResNet50 is more human-friendly to be observed than VGG19, as ResNet50 models have larger kernel sizes.
@@
In terms of Tiny ImageNet, we have Fig. 14 and Fig. 16 for the average of all images within 200 classes; also, Fig. 15 and Fig. 17 for the linear classifier visualization of all 200 classes in the dataset. We can observe the correlations between averaged images and the learned patterns show a similar phenomenon.
@@
Fig. 22 shows the performance gaps between ensembling on logits with greedy soups. From it, we can observe the similar phenomenon with the gaps of uniform soups, where the gaps grow when more models merged. But differently, the greedy soups break the going down trend as it can guarantee to maintain a relatively high accuracy level by only include the first model.
@@
We illustrate the ViT model (DeiT Tiny) on another dataset — Tiny ImageNet in Tab. 10. We can also see the more models are merged, the worse results from uniform soups can get. Also, we can observe the greedy soups, greedy ensemble on logits and greedy ensemble on features are not changed. It demonstrates the model-merging between 2 or more ViT models will not contribute to the performance. Thus, it shows the ViT models are not suitable for model-merging in the tested scenarios.
@@ 12 More Experimental Results of Model Magnitude Changes
The 2 models, 3 models, 5 models and 7 models accuracy vs. different magnification factors on Cifar100 are shown via Fig. 23, Fig. 24, Fig. 25 and Fig. 26, respectively.
@@
Similar to Cifar100, the 2 models, 3 models, 5 models, 7 models and 10 models accuracy vs. magnification factor on Cifar10 are shown via Fig. 27, Fig. 28, Fig. 29, Fig. 30 and Fig. 31, respectively. Moreover, we also visualize the uniform soup and logits ensemble accuracy vs. different magnification factors across different model numbers on the Cifar10 dataset in Fig. 32. We can observe similar phenomenons as discussed in the apepr. However, on the Cifar10, the gaps have different peak points than Cifar100, which is ×$\times 110$ (for Cifar100, it is ×$\times 100$).